# FSDP: Fast and Safe Data-Driven Overtaking Trajectory Planning for Head-to-Head Autonomous Racing Competitions

### v1

@@ III-A2 Data Selection
Second, the Range Filter ensures that data points are within a predefined range of values $[y_{\text{min}},y_{\text{max}}]$.
Third, the Confidence Interval Filter applies when $N_{\text{train}}>\frac{2}{3}N_{\text{target}}$, checking if the data points fall within the 95% confidence interval of the original SGP.
@@
6:$\mathcal{D}_{\text{filtered}}=\{\bm{x}_{j},y_{j}\in\mathcal{D}_{\text{incoming}}\mid y_{j}\in[y_{\text{min}},y_{\text{max}}]\}$
@@
12:$\mathcal{D}_{\text{merged}}\leftarrow\mathcal{D}_{\text{train}}\cup\mathcal{D}_{\text{filtered}}$
@@ III-B Efficient QP-Based Initial Trajectory Generation
[…]
where the constraints include position and velocity constraints at the start and end points.
@@ III-C Linearized Reference States Extraction Based on Differential Flatness
After the trajectory represented by the fifth-degree polynomial is obtained, the differential flatness property is used to find the linearized reference states, facilitating the lower-layer MPC optimization. For the kinematic bicycle model in Cartesian coordinates frame, the state vector $\mathbf{x_{c}}=[x,y,v,\theta]^{\top}$, the input vector $\mathbf{u}_{c}=[a_{t},\delta]^{\top}$ are defined. Where $[x,y]^{\top}$ is the position at the center of the rear wheel, $v$ is the longitudinal velocity, $a_{t}$ is the longitude acceleration, and $\delta$ is the steering angle. Based on the characteristics of differential flatness, all information is expressed through $x$ and $y$ as follows:
@@
where $L$ is the length of the wheelbase. Thus, the flat outputs $[x,y]^{\top}$ and their finite derivatives describe the arbitrary state and input information of the vehicle [17]. After selecting the reference points along the trajectory and coordinate transformation, the corresponding states $\mathbf{X}_{\text{ref}}$ and control inputs $\mathbf{U}_{\text{ref}}$ in the Frenet frame can be obtained using differential flatness. These reference points can then serve as the linearized reference points for MPC in the lower-layer planner.
@@ III-D MPC-Based Optimization for Kinematic Feasibility and Safety Guarantee
Assume the state sequence over the MPC prediction horizon $N$ is denoted by $\mathbf{X}=[\bm{x}_{1}^{\top},\dots,\bm{x}_{N}^{\top}]^{\top}\in\mathbb{R}^{N\cdot n_{x}}$, and $\mathbf{U}=[\bm{u}_{0}^{\top},\dots,\bm{u}_{N-1}^{\top}]^{\top}\in\mathbb{R}^{N\cdot n_{u}}$ denotes the control inputs sequence. In addition, we define an integrated state $\mathbf{X}^{=}[\mathbf{X}^{\top}-\mathbf{X}_{\text{ref}}^{\top},\mathbf{X}^{\top},\mathbf{X}^{\top}-\mathbf{X}_{\text{pre}}^{\top}]^{\top}$, where $\mathbf{X}_{\text{ref}}$ denotes the reference states sequence derived from differential flatness, and $\mathbf{X}_{\text{pre}}=[\bm{x}_{0}^{\top},\dots,\bm{x}_{N-1}^{\top}]^{\top}$. The optimization problem is formulated as:
@@ IV-A Experimental Setup
The evaluation will be divided into two parts: simulation and physical verification. The simulation scene uses the default F1TENTH simulator map, and the official F1TENTH simulator [3] is used to model the behavior of a 1:10 scale racing car in real-world scenarios. The CPU used for the experiments is an AMD Ryzen 7 5800H running at 3.2 GHz. Additionally, the physical experiment was conducted on an F1TENTH racecar based on [4], using an Intel NUC as the onboard computer. The composition of the car is shown in Fig. 3. The sensor suite includes a Hokuyo UST-10LX LiDAR with a maximum scan frequency of 40 Hz. Odometry is provided by the Vedder Electronic Speed Controller (VESC). Sensor fusion combines data from the LiDAR, the IMU embedded in the VESC, and odometry, utilizing the Cartographer to localize the vehicle and obtain its state.
During the experiment, both the ego car and the opponent car will follow the global minimum curvature racing line [18]. The trajectory tracking controller uses a tire-dynamics-based PP controller, with tire parameters identified through an on-track system identification method [19]. To better evaluate the overtaking ability of different algorithms, we define the speed-scaler $\mathcal{S}_{\text{max}}$ $=\frac{v_{\text{opp}}}{v_{\text{ego}}}\in[0,1]$. The ego vehicle consistently operates at its maximum speed $v_{\text{max}}$, while the opponent vehicle runs at a speed $v_{\text{opp}}$ lower than that of the ego vehicle. In the experiments, we evaluate different algorithms by progressively reducing the opponent’s speed until the ego vehicle is able to successfully complete the overtaking maneuver. The metrics used to evaluate the algorithm’s performance are as follows:
@@
$\mathcal{R}_{\text{ot/c}}$ is used to indicate the success rate of the overtaking algorithms. $l$ and $T$ represent the execution efficiency of the overtaking trajectory, while ${\bar{\dot{a}}}$ and $\bar{{\dot{\delta}}}$ describe the stability of the vehicle when tracking the overtaking trajectory.
Figure 3: Physical map and 1:10 scale autonomous F1TENTH racing platform.
@@ IV-B2 Overtake
As shown in Table II, the overtaking success rate of FSDP reached 90.90%, surpassing the performance of the listed SOTA. For the same $\mathcal{S}_{\text{max}}$, the overtaking success rate of FSDP is approximately 9.09% higher than that of the Predictive Spliner. The average speed during overtaking is $v_{o}=l/T$. For FSDP, its $v_{o}=4.67\,$\mathrm{m}\text{/}\mathrm{s}$$ the highest among SOTA methods, indicating that the trajectories it plans have higher execution efficiency. Due to the absence of dynamic opponent prediction, methods such as Frenet, GBO, and Spliner show lower overtaking success rates under lower values of $\mathcal{S}\text{max}$. Regarding trajectory quality, FSDP exhibits smaller values for ${\bar{\dot{a}}}$ and $\bar{{\dot{\delta}}}$ in simulation tests. Compared to the Predictive Spliner, these two metrics have decreased by 54.4% and 20.0%, respectively. This improvement can be attributed to the fact that FSDP simultaneously considers both the smoothness of the trajectory and the vehicle’s kinematic model. These results indicate that FSDP not only achieves high success rates but also ensures high performance in terms of trajectory quality. Fig. 6 visualizes the overtaking trajectories of all methods in simulation.
@@ IV-C2 Overtake
As shown in Table III, compared to SOTA, FSDP still demonstrates advantages in trajectory execution efficiency and smoothness. In comparison to the Predictive Spliner, the value of ${\bar{\dot{a}}}$ in FSDP decreased by 73.5%. With the highest $\mathcal{S}_{\text{max}}$, FSDP achieved an overtaking success rate of 71.43% on the real vehicle, representing the best performance among the SOTA methods. Compared to the Predictive Spliner, the success rate of FSDP is 8.93% higher, indicating improved safety. This result indicates that the trajectory planned by FSDP offers higher safety and smoothness on real racetracks. Fig. 9 visualizes the overtaking trajectories of all methods in the real vehicle experiments, and Fig. 10 shows FSDP’s overtaking screenshot.
@@ IV-D Compute
TABLE IV: Computational results of each overtaking strategy. Simulation is computed on the AMD Ryzen 7 CPU, and physical computations are done on the Intel NUC. The computational times are reported with their mean $\mu_{c}$ and standard deviation $\sigma_{c}$.

### v2

@@ III-A2 Data Selection
Second, the Range Filter ensures that data points are within a predefined range of values $y_{\text{min}},y_{\text{max}}]$ by removing physically implausible or sensor-faulty outliers to ensure data quality.
Third, the Confidence Interval Filter applies when $N_{\text{train}}>\frac{2}{3}N_{\text{target}}$, checking if the data points fall within the 95% confidence interval of the original SGP. This preserves data consistent with the current model’s predictions when the dataset is sufficiently large, preventing noise corruption and maintaining model stability.
@@
[…]
@@
[…]
@@ III-B Efficient QP-Based Initial Trajectory Generation
$\begin{split}&\min\,\,{\int_{0}^{T}\left\|P(t)-P^{\star}(t)\right\|^{2}dt}\\ &s.t.\quad\begin{array}[]{lc}P^{(j)}(T)=P^{\star(j)}(T)\\ P^{(j)}(0)=P^{\star(j)}(0),j\in\{0,1\}\end{array}\end{split}$ (9)
where the constraints include position and velocity constraints at the start and end points. The problem (9) is solved using OSQP [17].
@@ III-C Linearized Reference States Extraction Based on Differential Flatness
After the trajectory represented by the fifth-degree polynomial is obtained, the differential flatness property is used to find the linearized reference states, facilitating the lower-layer MPC optimization. For the kinematic bicycle model in Cartesian coordinates frame, the state vector $\mathbf{x}_{c}=[x,y,v,\theta]^{\top}$, the input vector $\mathbf{u}_{c}=[a_{t},\delta]^{\top}$ are defined. Where $x,y]^{\top}$ is the position at the center of the rear wheel, $v$ is the longitudinal velocity, $a_{t}$ is the longitude acceleration, and $\delta$ is the steering angle. Based on the characteristics of differential flatness, all information is expressed through $x$ and $y$ as follows:
@@
where $L$ is the length of the wheelbase. Thus, the flat outputs $x,y]^{\top}$ and their finite derivatives describe the arbitrary state and input information of the vehicle [18]. After selecting the reference points along the trajectory and coordinate transformation, the corresponding states $\mathbf{X}_{\text{ref}}$ and control inputs $\mathbf{U}_{\text{ref}}$ in the Frenet frame can be obtained using differential flatness. These reference points can then serve as the linearized reference points for MPC in the lower-layer planner.
@@ III-D MPC-Based Optimization for Kinematic Feasibility and Safety Guarantee
Assume the state sequence over the MPC prediction horizon $N$ is denoted by $\mathbf{X}=[\bm{x}_{1}^{\top},\dots,\bm{x}_{N}^{\top}]^{\top}\in\mathbb{R}^{N\cdot n_{x}}$, and $\mathbf{U}=[\bm{u}_{0}^{\top},\dots,\bm{u}_{N-1}^{\top}]^{\top}\in\mathbb{R}^{N\cdot n_{u}}$ denotes the control inputs sequence. In addition, we define an integrated state $\mathbf{X}^{*}=[\mathbf{X}^{\top}-\mathbf{X}_{\text{ref}}^{\top},\mathbf{X}^{\top},\mathbf{X}^{\top}-\mathbf{X}_{\text{pre}}^{\top}]^{\top}$, where $\mathbf{X}_{\text{ref}}$ denotes the reference states sequence derived from differential flatness, and $\mathbf{X}_{\text{pre}}=[\bm{x}_{0}^{\top},\dots,\bm{x}_{N-1}^{\top}]^{\top}$. The optimization problem is formulated as:
@@ IV-A Experimental Setup
The evaluation will be divided into two parts: simulation and physical verification. The simulation scene uses the default F1TENTH simulator map, and the official F1TENTH simulator [3] is used to model the behavior of a 1:10 scale racing car in real-world scenarios. The CPU used for the simulations is an AMD Ryzen 7 5800H running at 3.2 GHz. Additionally, the physical experiment was conducted on an F1TENTH racecar based on [4], using an Intel NUC as the onboard computer. The sensor suite includes a Hokuyo UST-10LX LiDAR with a maximum scan frequency of 40 Hz. Odometry is provided by the Vedder Electronic Speed Controller (VESC). Sensor fusion combines data from the LiDAR, the IMU embedded in the VESC, and odometry, utilizing the Cartographer to localize the vehicle and obtain its state.
During the experiment, both the ego car and the opponent car will follow the global minimum curvature racing line [19]. The trajectory tracking controller uses a tire-dynamics-based PP controller, with tire parameters identified through an on-track system identification method [20]. To better evaluate the overtaking ability of different algorithms, we define the speed-scaler $\mathcal{S}_{\text{max}}$ $\frac{v_{\text{opp}}}{v_{\text{ego}}}\in[0,1]$. The ego vehicle runs at its maximum trackable reference speed, while the opponent moves at the highest speed that still allows overtaking. The speed-scaler reflects their relative speed ratio and serves as an indicator of overtaking difficulty. In the experiments, we evaluate different algorithms by progressively reducing the opponent’s speed until the ego vehicle is able to successfully complete the overtaking maneuver. The metrics used to evaluate the algorithm’s performance are as follows:
@@
$\mathcal{R}_{\text{ot/c}}$ is used to indicate the success rate of the overtaking algorithms. $l$ and $T$ represent the execution efficiency of the overtaking trajectory, while $\bar{\dot{a}}$} and $\bar{{\dot{\delta}}}$ describe the stability of the vehicle when tracking the overtaking trajectory. The planner is only activated when an obstacle is detected; otherwise, it remains in racing mode [4].
@@ IV-B2 Overtake
As shown in Table II, the overtaking success rate of FSDP reached 90.90%, surpassing the performance of the listed SOTA. For the same $\mathcal{S}_{\text{max}}$, the overtaking success rate of FSDP is approximately 9.09% higher than that of the Predictive Spliner. The average speed during overtaking is $v_{o}=l/T$. For FSDP, its $v_{o}=4.67\,$\mathrm{m}\text{/}\mathrm{s}$\$ the highest among SOTA methods, indicating that the trajectories it plans have higher execution efficiency. Due to the absence of dynamic opponent prediction, methods such as Frenet, GBO, and Spliner show lower overtaking success rates under lower values of $\mathcal{S}_{\text{max}}$. Regarding trajectory quality, FSDP exhibits smaller values for $\bar{\dot{a}}$} and $\bar{{\dot{\delta}}}$ in simulation tests. Compared to the Predictive Spliner, these two metrics have decreased by 54.37% and 20.0%, respectively. This improvement can be attributed to the fact that FSDP simultaneously considers both the smoothness of the trajectory and the vehicle’s kinematic model. These results indicate that FSDP not only achieves high success rates but also ensures high performance in terms of trajectory quality. Fig. 5 visualizes the overtaking trajectories of all methods in simulation.
@@ IV-C2 Overtake
As shown in Table III, compared to SOTA, FSDP still demonstrates advantages in trajectory execution efficiency and smoothness. In comparison to the Predictive Spliner, the value of $\bar{\dot{a}}$} in FSDP decreased by 73.53%. With the highest $\mathcal{S}_{\text{max}}$, FSDP achieved an overtaking success rate of 71.43% on the real vehicle, representing the best performance among the SOTA methods. Compared to the Predictive Spliner, the success rate of FSDP is 8.93% higher, indicating improved safety. This result indicates that the trajectory planned by FSDP offers higher safety and smoothness on real racetracks. Fig. 8 visualizes the overtaking trajectories of all methods in the real vehicle experiments, and Fig. 9 shows FSDP’s overtaking screenshot.
@@ IV-D Compute
TABLE IV: Computational results of each overtaking strategy. The computational times are reported with their mean $\mu_{c}$ and standard deviation $\sigma_{c}$.